OpenStudy (anonymous):

How do I factor this: 16k²-1

OpenStudy (anonymous):

(4k + 1) (4k - 1)

OpenStudy (saifoo.khan):

Hav u tried to solve?

OpenStudy (anonymous):

saifoo, stop trying to make people learn, jeez :p

OpenStudy (anonymous):

\[x^2-y^2=(x+y)(x-y)\]

OpenStudy (saifoo.khan):

LOl, the accounts of both asker's and answer's can get banned due to that.

OpenStudy (saifoo.khan):

This is marked as Cheating.

OpenStudy (anonymous):

yes i agree

OpenStudy (anonymous):

\[16k²-1\] just try it first , bbut i am giving u solution . \[(4k)^2-(1)^2\] as we know that \[a^2-b^2=(a+b)(a-b)\] hence here \[(4k)^2-(1)^2=(4k+1)(4k-1)\]

OpenStudy (anonymous):

thank you, I wasn't sure how to factor this specific equation and I was just stumped

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

difference of squares